# Deep Learning for Retinal Degeneration Assessment: A Comprehensive Analysis of the MARIO AMD Progression Challenge

### v1

@@ 5.1 Summary of methods proposed for each team
DF41 OCTIP* ResNet50,ViTLarge-MAE (PPMAE) CE Ensembling, (average) RandomHorizontalFlip, RandomVerticalFlip,RandomRotation, ColorJitter, RandomPerspective, GaussianBlur N N N N &\&& Y/ PPMAE https://github.com/pzhangwj/mario_challenge_code
@@ 5.1.3 Cemrg - Team Summary
An ablation study was performed by comparing the performance of $MobileNetV3\_base$ (without ensembling or TTA) against $MobileNetV3\_ensemble$ (with ensembling and TTA).
@@
Ensemble outputs were combined through averaging or a weighted sum. An ablation study comparing $MobileNetV3\_base$ and $MobileNetV3\_ensemble$ models was conducted to highlight the benefits of ensembling and TTA.
@@ 5.1.6 FERLIV - Team Summary
The FERLIV team proposed a modular, fully transformer-based approach for monitoring the progression of wet AMD using two registered OCT B-scans from two consecutive medical exams. The method consists of three transformer encoder architectures: Feature Encoder, Change Encoder, and Diagnosis Encoder. The Feature Encoder is the first part of the method, where a large Vision Transformer [51] encodes a pair of OCT B-scans into two sets of local feature vectors. The Change Encoder detects visual changes in a retina using a dual Multi-Change Captioning Transformer [52] with a co-attention mechanism [53]. The outputs of the Change Encoder are concatenated based on their corresponding 2D positions and forwarded to the third and final transformer encoder, i.e., the Diagnosis Encoder, which quantifies changes in disease progression. The Diagnosis Encoder uses a self-attention mechanism and is followed by the classification head. During preprocessing, OCT B-scans are resized and cropped to $224\times 224$ pixels, followed by normalization. Before training, the Feature Encoder is initialized with weights from the pretext task, where RETFound’s Vit-Large model [54] is used as a Segmenter encoder [55] for the segmentation of retinal layers on the OCT MS and Healthy Control dataset [56]. This public dataset is acquired on the same device as the MARIO dataset, addressing variability caused by inter-device transfer. Pretext task results in improved performance for Task 1, as changes in retinal structures are crucial for monitoring disease progression. All three encoders are trained using the AdamW optimizer with cosine learning rate scheduling and weight decay. A weighted categorical cross-entropy loss function is used to address the class imbalance. Augmentation techniques include random resized cropping and random horizontal flipping to improve generalization.
@@ 6.1.2 Final Ranking Computation
Final $\text{Final Score}=0.35\times f_{\text{task\_1}}+0.65\times f_{\text{task\_2}}$ (3)
@@ 6.4.1 Task 1
Figure 14: Heatmap of accuracy as a function of delta time for Task 1 on the Tlemcen dataset. The horizontal axis represents the delta time (tt) between baseline and follow-up scans, while the vertical axis lists the competing teams. Color intensity corresponds to the accuracy, with brighter areas indicating higher predictive performance. This visualization provides an in-depth temporal analysis of model generalization across varying time intervals.
@@ 6.4.3 Ranking on the test set from Algeria
team $F1\_$score Rk-correlation Specificity final score Task 1
@@ 6.7.1 Ranking boostraping
To estimate the stability of the rankings, we employed a bootstrap resampling procedure, following the methodology proposed by Wiesenfarth et al. [57]. This involves resampling predictions with replacement and recalculating performance metrics over multiple iterations to generate a distribution of rankings for each team.
@@ 6.7.2 McNemar Test for Model Comparisons
Figure 22: McNemar test results for pairwise comparisons of team predictions on both tasks in Brest (top row) and Tlemcen (bottom row). Each cell represents $-\log_{10}(p)$ for a McNemar test between two models. Only statistically significant values (i.e., $p<0.01$) are shown in the upper triangle of the matrix. Asterisks denote significance levels (* for $p<0.001$, ** for $p<0.0001$). These heatmaps highlight which teams produce significantly different predictions, with denser significance patterns observed in Task 2 across both cities.

### v3

@@ 5.1 Summary of methods proposed for each team
DF41 OCTIP* ResNet50,ViTLarge-MAE (PPMAE) CE Ensembling, (average) RandomHorizontalFlip, RandomVerticalFlip,RandomRotation, ColorJitter, RandomPerspective, GaussianBlur N N N N &\& Y/ PPMAE https://github.com/pzhangwj/mario_challenge_code
@@ 5.1.3 Cemrg - Team Summary
An ablation study was performed by comparing the performance of MobileNetV3_baseMobileNetV3\_base (without ensembling or TTA) against MobileNetV3_ensembleMobileNetV3\_ensemble (with ensembling and TTA).
@@
Ensemble outputs were combined through averaging or a weighted sum. An ablation study comparing MobileNetV3_baseMobileNetV3\_base and MobileNetV3_ensembleMobileNetV3\_ensemble models was conducted to highlight the benefits of ensembling and TTA.
@@ 5.1.6 FERLIV - Team Summary
The FERLIV team proposed a modular, fully transformer-based approach for monitoring the progression of wet AMD using two registered OCT B-scans from two consecutive medical exams. The method consists of three transformer encoder architectures: Feature Encoder, Change Encoder, and Diagnosis Encoder. The Feature Encoder is the first part of the method, where a large Vision Transformer [50650] encodes a pair of OCT B-scans into two sets of local feature vectors. The Change Encoder detects visual changes in a retina using a dual Multi-Change Captioning Transformer [qiu2021describing] with a co-attention mechanism [lu2019vilbert]. The outputs of the Change Encoder are concatenated based on their corresponding 2D positions and forwarded to the third and final transformer encoder, i.e., the Diagnosis Encoder, which quantifies changes in disease progression. The Diagnosis Encoder uses a self-attention mechanism and is followed by the classification head. During preprocessing, OCT B-scans are resized and cropped to $224\times 224$ pixels, followed by normalization. Before training, the Feature Encoder is initialized with weights from the pretext task, where RETFound’s Vit-Large model [zhou2023foundation] is used as a Segmenter encoder [strudel2021segmenter] for the segmentation of retinal layers on the OCT MS and Healthy Control dataset [he2019retinal]. This public dataset is acquired on the same device as the MARIO dataset, addressing variability caused by inter-device transfer. Pretext task results in improved performance for Task 1, as changes in retinal structures are crucial for monitoring disease progression. All three encoders are trained using the AdamW optimizer with cosine learning rate scheduling and weight decay. A weighted categorical cross-entropy loss function is used to address the class imbalance. Augmentation techniques include random resized cropping and random horizontal flipping to improve generalization.
@@ 6.1.2 Final Ranking Computation
$\text{Final Score}=0.35\times f_{\text{task\_1}}+0.65\times f_{\text{task\_2}}$ (3)
@@ 6.4.1 Task 1
Figure 14: Heatmap of accuracy as a function of delta time for Task 1 on the Tlemcen dataset. The horizontal axis represents the delta time ($t$) between baseline and follow-up scans, while the vertical axis lists the competing teams. Color intensity corresponds to the accuracy, with brighter areas indicating higher predictive performance. This visualization provides an in-depth temporal analysis of model generalization across varying time intervals.
@@ 6.4.3 Ranking on the test set from Algeria
team F1_F1\_score Rk-correlation Specificity final score Task 1
@@ 6.7.1 Ranking boostraping
To estimate the stability of the rankings, we employed a bootstrap resampling procedure, following the methodology proposed by Wiesenfarth et al. [Wiesenfarth2021]. This involves resampling predictions with replacement and recalculating performance metrics over multiple iterations to generate a distribution of rankings for each team.
@@ 6.7.2 McNemar Test for Model Comparisons
Figure 22: McNemar test results for pairwise comparisons of team predictions on both tasks in Brest (top row) and Tlemcen (bottom row). Each cell represents $\log_{10}(p)$ for a McNemar test between two models. Only statistically significant values (i.e., $p<0.01$) are shown in the upper triangle of the matrix. Asterisks denote significance levels (* for $p<0.001$, ** for $p<0.0001$). These heatmaps highlight which teams produce significantly different predictions, with denser significance patterns observed in Task 2 across both cities.